Differentiate the following functions with respect to x: (i) ln(1 + sin2 x) (ii) xx.
(c) Evaluate the integral 2x3 −4x−8 /x4 −x3 +4x2 −4x dx.
1
Expert's answer
2020-06-29T17:41:08-0400

1.i. "ln [ 1 + sin 2x]"

Differentiating with respect to "x" :

"d( ln [ 1 + sin2x])\/dx"

"=[2 cos(2x)] \/ [ 1 + sin 2x]" (Answer)

ii. "x^x"

Let A = "x^x"

Taking natural logarithm on both sides:

"ln A= x ln x"

Differentiating with respect to x:

"d[ ln A]\/dx = d[x ln x] \/ dx"

or, "1 \/ A * d[A]\/dx = x\/x + ln x"

or, "dA\/dx = (1 + ln x) * A"

Putting A = "x^x"

"dA\/dx = [1 + ln x] * X^x Answer"


c. Solve "\int ( 2x^3 - 4x - 8/x^4 - x^3 + 4x^2 -4x) dx"

= "2x^4\/4 - 4x^2\/2 - 8x^{-3} \/( -3) - x^4\/4 + 4x^3\/3 - 4x^2\/2 + c"

where C is an integration constant

= "x^4\/2 - 2x^2 + 8(x^-3 \/ 3) - x^4\/4 + 4x^3\/3 - 2x^2 + c"

="[x^4\/2 - x^4\/4] - [ 2x^2 + 2 x^2] + 8x^-3\/3 + 4x^3\/3 + c"

="[x^4\/4] - 4x^2 + 8\/3x^3 + 4x^3\/3 + c [Answer]"